Perfect Gases

0.675kg of gas at 1.4 MN/m2 and 2800C is expanded to four times the original volume according to the law PV1.3 =C Determine:

(a) the original and final volume of the gas,

(b) the final pressure of the gas

(c) the final temperature of the gas

Take R=0.287Kj/kg K



1
Expert's answer
2022-03-10T10:49:20-0500

m1 = 0.675 kg

p1 = 1.4 MN/m2

T1 = 280°C

 R = 0.287 kJ/kg K,

v2 = 4 V1


p1v1 = mRT1


(a.) v1 = mRT1 / T1


v1= "\frac{0.675\times0.287\times553.15}{14\times10^3}" = 0.0077


V2= 4V1= 4"\times"10-3 = 0.03m3



(b.) P1V1n = P2V2n


P2= P1["\frac{V_1}{V_2}"]n= 14"\times" 103kPa["\frac{1}{4}"]1.3 = 2.3"\times"103 kPa



(c.) "\frac{P_1V_1}{T_1}= \frac{P_2V_2}{T_2}"


T2 = "\frac{P_2}{P_1}\times\frac{V_2}{V_1}\times(T_1)"


T2 = "\frac{2.3\times10^3}{14\times10^3}\times\frac{4}{1}\times553.15"


T2 = 363.499 - 280 = 83.5°C


T2 = 83.5°C